If the perimeter is 46 centimeters. And the length is 3cm more than 3 times the length of the width. What is the length and the width for this rectangle?
please respond ASAP!

Answer by mbmodak(32)

Let width of rect. Be x
length=3x+3
Perimeter of rect.=2(length+breadth)
46=2[x+(3x+3)]
46=2x+6x+6
46=8x+6
46-6=8x
40=8x
x=40/8=5
breadth=5 cm & length=3*5+3=18
46=2[5+(5*3+3)
46=2[5+18]
46=2*23